OpenStudy (mysticfury):

Use elimination to find the solution to the system of equations. -x + 2y = 15 5x + 5y = -30 A. x = 3, y = 9 B. x= -5, y = 5 C. x = -9, y = 3 D. x = -3, y = -3

OpenStudy (welshfella):

first multiply the first equation by 5 then add this to the second equation this will eliminate the 'x' terms.

OpenStudy (mysticfury):

what do you mean by multiply it by 5?

OpenStudy (welshfella):

multiply every term in the equation by 5 -x + 2y = 15 * by 5 gives -5x + 10y = 75

OpenStudy (welshfella):

now adding this to second equation eliminates 'x' -5x + 10y = 75 + 5x + 5y = -30

OpenStudy (welshfella):

can you do the addition? @MysticFury